OpenStudy (anonymous):

If xy+x+y+1=0,x+ay-3=0 are concurrent,then a=?

OpenStudy (kc_kennylau):

https://www.desmos.com/calculator/5zqq4gg9ck

OpenStudy (phi):

strange equation ! xy+x+y+1=0 if we factor it this way: x(y+1) + (y+1) = 0 (x+1)(y+1) = 0 we get either x+1=0 or y+1=0 x=-1, or y=-1 this looks like two lines that intersect at -1,-1

OpenStudy (kc_kennylau):

I think it needs 3 lines to be concurrent?

OpenStudy (phi):

so I think they want to know what "a" makes x+ay-3=0 go through the point -1,-1

OpenStudy (phi):

? -1+a(-1) - 3 =0 a= -4

OpenStudy (anonymous):

i think there are a lot of 'a' make xy+x+y+1=0,x+ay-3=0 concurrent ;